OpenStudy (anonymous):

Simplify: (3x + 12) (3x - 12).

OpenStudy (anonymous):

would that give me the answer

OpenStudy (texaschic101):

Are you familiar with the FOIL method ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I do not understand it please help me!!!!

OpenStudy (anonymous):

guys pleaaase !

OpenStudy (texaschic101):

F -first O - outside I - inside L - last (3x + 12)(3x - 12) F - multiply first terms in each set -- 3x * 3x = 9x^2 O - multiply outside terms -- 3x * -12 = -36x I - multiply inside terms -- 12 * 3x = 36x L - multiply last terms -- 12 * -12 = -144 now put it all together 9x^2 - 36x + 36x - 144 combine like terms 9x^2 - 144 (the x's cancel each other out) Do you understand ?

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

yes I understand thank you!!!! how about this? Solve: 2(x - 3) + 2(x - 7) = 40

OpenStudy (texaschic101):

2(x - 3) + 2(x - 7) = 40 2x - 6 + 2x - 14 = 40 4x - 20 = 40 4x = 40 + 20 4x = 60 x = 60/4 = 15

OpenStudy (anonymous):

that answer does not appear here the answers that appear are:x = 12.5 x = 15 x = 5 no solution infinitely many solutions

OpenStudy (texaschic101):

x = 15

OpenStudy (anonymous):

oooooo ok youre really smart thank you Simplify: (x + 9) (x + 8). Answer x2 + 17x + 72 x2 + 72x + 17 x2 + 17 x2 + 72

OpenStudy (anonymous):

im having a really hard time with algebra!

OpenStudy (texaschic101):

Thats the FOIL method (x + 9)(x + 8) x^2 + 8x + 9x + 72 x^2 + 17x + 72

OpenStudy (anonymous):

is this the foil method too?? Simplify: (4b + 3) (2b - 5). Answer 8b2 - 14b - 15 8b2 + 9b - 15 8b2 + 6b + 15 8b2 - 15

OpenStudy (texaschic101):

yes it is the FOIL method...you wanna try it...I will walk you through it

OpenStudy (anonymous):

Simplify: (3x - 8) (2x - 2). Answer 6x2 + 16 6x2 - 22x + 16 6x2 - 16x + 16 6x2 - 6x - 16

OpenStudy (anonymous):

Simplify: (x - 1)(x2 + 10x + 4).

OpenStudy (anonymous):

wat?

OpenStudy (texaschic101):

x^3 + 9x^2 - 6x - 4

OpenStudy (texaschic101):

the (3x - 8) (2x - 2) is the FOIL method

OpenStudy (anonymous):

Simplify: (4b + 3) (2b - 5). Answer 8b2 - 14b - 15 8b2 + 9b - 15 8b2 + 6b + 15 8b2 - 15

OpenStudy (anonymous):

Simplify: (3x - 8) (2x - 2). Answer 6x2 + 16 6x2 - 22x + 16 6x2 - 16x + 16 6x2 - 6x - 16

OpenStudy (texaschic101):

lets do this one together... (4b + 3)(2b - 5) First -- 4b * 2b = ?

OpenStudy (anonymous):

5

OpenStudy (texaschic101):

no.....4b * 2b ? do you know how to do this

OpenStudy (texaschic101):

whats 4 * 2 and b * b ?

OpenStudy (anonymous):

I don't know!!

OpenStudy (anonymous):

i need the answers i don't understand this!

OpenStudy (texaschic101):

(4b + 3)(2b - 5) 8b^2 - 20b + 6b - 15 8b^2 - 14b - 15

OpenStudy (texaschic101):

(3x - 8)(2x - 2) 6x^2 - 6x - 16x + 16 6x^2 - 22x + 16

OpenStudy (anonymous):

how about Simplify: (3x - 8) (2x - 2). Answer 6x2 + 16 6x2 - 22x + 16 6x2 - 16x + 16 6x2 - 6x - 16

OpenStudy (texaschic101):

just did it

OpenStudy (anonymous):

thank you how about Simplify: (3x - 8) (2x - 2). Answer 6x2 + 16 6x2 - 22x + 16 6x2 - 16x + 16 6x2 - 6x - 16

OpenStudy (texaschic101):

I already did that one...check above

OpenStudy (anonymous):

Simplify: (x - 1)(x2 + 10x + 4). Answer x3 + 10x2 + 4x - 1 x3 + 9x2 - 6x - 4 x3 + 10x2 + 4x - 4 x3 + 9x2 - 6x - 1

OpenStudy (texaschic101):

x^3 + 9x^2 - 6x - 4

OpenStudy (anonymous):

Multiply: 5xy(2x - y) Answer 10xy - 5xy 10x2y - 5xy2 10x2 - 5xy + 25y2 10x2 - xy + y2

OpenStudy (texaschic101):

5xy(2x - y) 10x^2y - 5xy^2

OpenStudy (anonymous):

Simplify: 2x(4x2 - 3x - 9) Answer 8x2 - 3x - 9 8x2 - 6x - 18 8x3 - 6x2 - 18x 8x3 - 3x2 - 9x

OpenStudy (texaschic101):

I am sorry but I have to go now...good luck

OpenStudy (anonymous):

thank you !!!!

OpenStudy (texaschic101):

your welcome